# Solving Oscillator Ordinary Differential Equations via Soft-constrained Physics-informed Neural Network with Small Data

### v2

@@ 2.1.1 Governing Equations
Consider a physical system defined on a spatial or spatio-temporal domain $\Omega\subseteq\mathbb{R}^{d}$, where the unknown $u(\bm{x}):\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ is system state variable that are functions of spatial or temporal coordinates $\bm{x}\in\Omega$. For time-independent systems, $\bm{x}=(x_{1},\ldots,x_{d})$; for time-dependent systems, $\bm{x}=(x_{1},\ldots,x_{d-1},t)$. The physical laws of the dominant system are often characterized in terms of ODEs/PDEs, and these equations are known as the governing equations [HLZ+23], the general form of which is given by
@@
For time-dependent systems (i.e., dynamical systems), initial conditions $\mathcal{I}(u;\theta)(x,t_{0})=0,x\in\Omega_{0}$ (Eq. (2)) need to be set for the state variables (and sometimes their derivatives) at the initial moment $t_{0}$. $\theta\in\Theta$ parameterizes or controls the system, where $\theta$ can be a vector or a function included in the control equation. For systems characterized by PDEs, we also need to constrain the state variables on the boundaries of the spatial domain $\partial\Omega$ in order to make the system well-posed. For the boundary points $x\in\partial\Omega$, we have boundary conditions $\mathcal{B}(u;\theta)(x,t)=0,x\in\partial\Omega$ (Eq. (3)). If there are no constraints on the initial and boundary conditions, then $\mathcal{I}(u;\theta)\triangleq 0$ and $\mathcal{B}(u;\theta)\triangleq 0$ [HLZ+23].
@@ 2.1.2 Framework of PINN via Soft Constrains
As shown in Figure 1, suppose there is a system obeying Eq. (1) and a dataset $\{u(\bm{x}_{i})\}_{i=1,\ldots N}$. It is then possible to construct the neural network $u_{w}(\bm{x})$ and train it with the following loss function, namely
$\begin{split}\mathcal{L}_{total}=\mathcal{L}_{data}+\underbrace{\mathcal{L}_{gov}+\mathcal{L}_{IC}+\mathcal{L}_{BC}+\cdots}_{\mathcal{L}_{phys}}\\ =\frac{\lambda_{d}}{N}\sum_{i=1}^{N}\|u_{w}(\bm{x}_{i})-u(\bm{x}_{i})\|^{2}+\frac{\lambda_{g}}{|\Omega|}\int_{\Omega}\|\mathcal{F}(u_{w};\theta)(\bm{x})\|^{2}\mathrm{d}\bm{x}+\\ \frac{\lambda_{i}}{|\Omega_{0}|}\int_{\Omega_{0}}\|\mathcal{I}(u_{w};\theta)(\bm{x})\|^{2}\mathrm{d}\bm{x}+\frac{\lambda_{b}}{|\partial\Omega|}\int_{\partial\Omega}\|\mathcal{B}(u_{w};\theta)(\bm{x})\|^{2}\mathrm{d}\bm{x}+\cdots.\end{split}$ (4)
Where, $\mathcal{L}_{data}=\frac{1}{N}\sum_{i=1}^{N}\|u_{w}(\bm{x}_{i})-u(\bm{x}_{i})\|^{2}$ is the regular data loss of PINN in matching the labled datasets; $\mathcal{L}_{gov}=\int_{\Omega}\|\mathcal{F}(u_{w};\theta)(\bm{x})\|^{2}\mathrm{d}\bm{x}$ is the residual loss that makes the network $u_{w}$ satisfy the governing equation constraints; $\mathcal{L}_{IC}=\int_{\Omega_{0}}\|\mathcal{I}(u_{w};\theta)(\bm{x})\|^{2}\mathrm{d}\bm{x}$ and $\mathcal{L}_{BC}=\int_{\partial\Omega}\|\mathcal{B}(u_{w};\theta)(\bm{x})\|^{2}\mathrm{d}\bm{x}$ are the corresponding losses that make PINN satisfy the initial and boundary conditions, respectively.
@@
In order to compute (4), several integral terms need to be evaluated, which involve the computation of higher-order derivatives of $u_{w}(\bm{x})$. PINN utilizes automatic differentiation of computational maps to calculate these derivative terms. The integrals are then approximated using a set of collocation points, which can be sampled using the Monte-Carlo method. We use $N_{d}$, $N_{g}$, $N_{i}$, $N_{b}$ to denote the number of corresponding data or collocation points. Then, the loss function can be approximated as [HLZ+23]
$\begin{split}\mathcal{L}_{total}=\frac{\lambda_{d}}{N_{d}}\sum_{i=1}^{N}\|u_{w}(\bm{x}_{i})-u(\bm{x}_{i})\|^{2}+\frac{\lambda_{g}}{N_{g}}\sum_{i=1}^{N_{g}}\|\mathcal{F}(u_{w};\theta)(\bm{x}_{i})\|^{2}+\\ \frac{\lambda_{i}}{N_{i}}\sum_{i=1}^{N_{i}}\|\mathcal{I}(u_{w};\theta)(\bm{x}_{i})\|^{2}+\frac{\lambda_{b}}{N_{b}}\sum_{i=1}^{N_{b}}\|\mathcal{B}(u_{w};\theta)(\bm{x}_{i})\|^{2}+\cdots.\end{split}$ (5)
@@ 2.1.3 Implementation Based on DeepXDE
We implemented soft-constrained PINN with DeepXDE [LMMK21] and compared the results with a PyTorch-based implementation [BB23] (see  2.2.2 for details). DeepXDE is well-structured and highly configurable. Code written in DeepXDE is shorter, closer to mathematical formulas, and more computationally efficient. Solving differential equations in DeepXDE is akin to “building blocks” using built-in modules that specify the computational domain (geometric and temporal), differential equations, boundary/initial conditions, constraints, training datasets, neural network structure, optimization algorithms, and hyper-parameters, etc., all of which are loosely coupled.
@@ 2.2 Results of Soft-constrained PINN Solving Oscillator ODEs
For direct comparison, we choose to solve the same linear and nonlinear oscillator ODEs (i.e., Tutorial Example and Van der Pol osillators) as in  [BB23].
@@ 2.2.1 PINN vs NN to Solving Linear Oscillator ODE
The only difference between solving with the soft-constrained PINN approach in this example and a conventional NN is the addition of $\mathcal{L}_{phys}$ to the total loss function. Since the initial condition in the conventional NN is already included in the training set (i.e., the first training point on the left), only $\mathcal{L}_{gov}$ satisfying Eq. (6) is added to total loss. This is accomplished by adding extra collocation points (see Eq. (5) ), in this example sampling at least 50 collocation points uniformly throughout the entire time domain. It could be understood that the training points are supervised and the collocation points are unsupervised. The exact number of collocation points required depends on the number of neural network parameters, i.e. the size of the problem being solved. The distribution of collocation points (uniform/random) also has an impact on the results [WZT+23]. The learning rate and loss weights also affect the convergence of the gradient descent algorithm [BB23]. Here the loss weights were taken as $\lambda_{d}=1.0,\lambda_{r}=6\times 10^{-2}$ through hyperparameter optimization. The loss weights act as a balance between the dual drivers of data and physical information.
@@
Figure 2: Result Comparison of PINN vs NN Solving Linear Oscillator–Case 1. ($N_{data}$–No. of training data, $N_{c}$–No. of collocation points, all the same below).
@@
Figure 3: Result Comparison of PINN vs NN Solving Linear Oscillator–Case 2.
Minimalistic Training Data. So, what is the minimum amount of training data required for the PINN method? For this example, if only one point, the initial value, is taken and the collocation points are kept to be 50, a good solution is obtained, as shown in Fig. 4(a), and the loss function descent process is shown in Fig. 4(b). [BB23] also show that when the nonlinearity of the problem is weak, a very small amount of labeled training data plus a few collocation points is sufficient to predict the solution. In the minimalist case, just as classical analytic or numerical integration methods require only one and two initial conditions for solving first- and second-order differential equations, respectively. PINN also require only one training point (initial value) and two training points containing the initial value point, respectively (see Fig. 7(b)).
@@
Figure 4: Minimalistic Training Data Example of First-order Linear ODE.
Accelerating Training Time. Compared to NN, PINN can significantly shorten the training time because it requires only a small amount of data (minimal labeled training data plus a small number of collocation points). On the other hand, as its loss function becomes more complex with the addition of $\mathcal{L}_{phys}$, it requires more computations and sometimes even more iteration epoches when training, which in turn drags down the training efficiency. We performed a validation test of Primer oscillator on DeepXDE 1.11.1 based on a CPU platform (Windows10, Intel Core i9-9900K @3.6GHz, 32GB Memory). The training time is the average of 5 training sessions (24,000 epoches per session, usually 24,000 epoches are not needed before convergence, and 24,000 epoches are trained for fair comparison), and the results are shown in Fig. 5. It can be seen that, for linear oscillator ODEs, the training is accelerated when the gains obtained along with the reduction in the amount of data outweigh the delay caused by the increase in the loss function terms. It should be emphasized that since Primer oscillator is a simple linear ODE, we used the same base network ($3\times 32$) for ease of comparison. PINN can accelerate the training even further by reducing the number of network parameters, because PINN need less network parameters than NN due to its small data characteristics. For example, for Fig. 5(b), if the base network is reduced to $3\times 16$, it also converges well while the training time is reduced to 15.30s.
@@
Figure 5: Training Time of PINN vs NN Solving Primer Oscillator via DeepXDE: (a) 16.56s by NN with Min. 50 training points; (b) 15.23s by PINN with Min. 1 initial value training point + 50 collocation points.
@@ 2.2.2 PINN Solving Non-linear Oscillator ODE
The soft-constrained PINN approach also performs well in solving nonlinear oscillator ODEs [BB23]. We successfully solved the typical strongly nonlinear Van der Pol (VDP) oscillator based on DeepXDE, and demonstrated the tolerance of PINN to noisy data as well as the light-code and training speed advantages of the DeepXDE implementation of PINN. The ODE of VDP is
@@
Noisy Data Tolerated. The PINN method is used to solve the nonlinear VDP oscillator with small and noisy data. As with solving the Primer oscillator, the fully connected base network of $3\times 32$ was still used, which reflects the strong expressive power of neural networks inherited by PINN. The relevant parameter settings (i.e., $\epsilon$, number of training data points, number of collocation points, normal noise variance) are illustrated in Fig. 6, and other hyperparameter settings can be found in the code. Exact solution was generated by the 4th-order Runge-Kutta method. As can be seen from Fig. 6: 1) With the same amount of training data and collocation points, the results obtained from training with noisy data (right column) are basically identical to the noiseless results (left column), showing the good inclusiveness of the PINN method against noise. 2) As the degree of nonlinearity increases (by increasing $\epsilon$), more training and collocation points are needed to be arranged in regions with significant nonlinear features to obtain an exact solution. When $\epsilon=1,3,5$, the least number of training and collocation points are $N_{data}=28,32,38;N_{c}=15,25,40$, respectively. And 3) as the nonlinearity increases, the tolerance to noise decreases: when $\epsilon$ increases from 1 to 5, the $\sigma$ that can be tolerated decreases from 0.1 to 0.05; even increasing the number of data points has a limited effect on improving noise tolerance.
@@
PINN Implementation Code via DeepXDE vs PyTorch. Comparison of Code  1 and Code  2 shows that the overall process (seen in Fig. 1) of realizing soft-constrained PINN based on DeepXDE or PyTorch is consistent. DeepXDE has a better customized encapsulation of the PINN method and thus requires less coding. With the same setup for solving VDP, the number of code lines (about 40 lines) via DeepXDE implementation is roughly one-third as the PyTorch implementation. In addition, the training efficiency of DeepXDE is significantly better than that of PyTorch, thanks to the optimization of DeepXDE for PINN [LMMK21]. For the example in Figure 5(b), PyTorch-based training took 43.10s with the same settings, which is about 1.8 times more than that of DeepXDE.
@@ 48 Loss_total = Loss_data + Loss_phys
PINN Computational Efficiency @CPU vs @GPU. The PINN method does not depend on big data so it can be efficiently computed on platforms such as CPU or GPU. We compared the training time for solving Primer, VDP oscillators based on DeepXDE on CPU and GPU (CPU as before, GPU is Nvidia GTX1080Ti) respectively, taking the average of 5 training sessions (24000 epoches each session). Base networks use $2\times 32$ and $3\times 32$ fully connected hidden layers, respectively. As shown in Table  2, the training acceleration effect of GPU on PINN solving Primer and VDP is obvious: the training duration is reduced from 16.02s and 17.60s to 8.73s and 13.15s, respectively. When the base network is reduced to $3\times 16$, the training duration results show a consistent trend.
Table 2: PINN Training Time @CPU vs @GPU.
Training Time (second) @CPU, DeepXDE @GPU, DeepXDE

### v4

@@ 2.1.1 Governing Equations
Consider a physical system defined on a spatial or spatio-temporal domain $\Omega\subseteq\mathbb{R}^{d}$, where the unknown $u(\boldsymbol{x}):\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ is system state variable that are functions of spatial or temporal coordinates $\boldsymbol{x}\in\Omega$. For time-independent systems, $\boldsymbol{x}=(x_{1},\ldots,x_{d})$; for time-dependent systems, $\boldsymbol{x}=(x_{1},\ldots,x_{d-1},t)$. The physical laws of the dominant system are often characterized in terms of ODEs/PDEs, and these equations are known as the governing equations [33], the general form of which is given by
@@
For time-dependent systems (i.e., dynamical systems), initial conditions $\mathcal{I}(u;\theta)(x,t_{0})=0,x\in\Omega_{0}$ (Eq. (2)) need to be set for the state variables (and sometimes their derivatives) at the initial moment $t_{0}$. $\theta\in\Theta$ parameterizes or controls the system, where $\theta$ can be a vector or a function included in the control equation. For systems characterized by PDEs, we also need to constrain the state variables on the boundaries of the spatial domain $\partial\Omega$ in order to make the system well-posed. For the boundary points $x\in\partial\Omega$, we have boundary conditions $\mathcal{B}(u;\theta)(x,t)=0,x\in\partial\Omega$ (Eq. (3)). If there are no constraints on the initial and boundary conditions, then $\mathcal{I}(u;\theta)\triangleq 0$ and $\mathcal{B}(u;\theta)\triangleq 0$ [33].
@@ 2.1.2 Framework of PINN via Soft Constrains
As shown in Figure 1, suppose there is a system obeying Eq. (1) and a dataset $\{u(\boldsymbol{x}_{i})\}_{i=1,\ldots N}$. It is then possible to construct the neural network $u_{w}(\boldsymbol{x})$ and train it with the following loss function, namely
$\begin{split}\mathcal{L}_{total}=\mathcal{L}_{data}+\underbrace{\mathcal{L}_{gov}+\mathcal{L}_{IC}+\mathcal{L}_{BC}+\cdots}_{\mathcal{L}_{phys}}\\ =\frac{\lambda_{d}}{N}\sum_{i=1}^{N}\|u_{w}(\boldsymbol{x}_{i})-u(\boldsymbol{x}_{i})\|^{2}+\frac{\lambda_{g}}{|\Omega|}\int_{\Omega}\|\mathcal{F}(u_{w};\theta)(\boldsymbol{x})\|^{2}\mathrm{d}\boldsymbol{x}+\\ \frac{\lambda_{i}}{|\Omega_{0}|}\int_{\Omega_{0}}\|\mathcal{I}(u_{w};\theta)(\boldsymbol{x})\|^{2}\mathrm{d}\boldsymbol{x}+\frac{\lambda_{b}}{|\partial\Omega|}\int_{\partial\Omega}\|\mathcal{B}(u_{w};\theta)(\boldsymbol{x})\|^{2}\mathrm{d}\boldsymbol{x}+\cdots.\end{split}$ (4)
Where, $\mathcal{L}_{data}=\frac{1}{N}\sum_{i=1}^{N}\|u_{w}(\boldsymbol{x}_{i})-u(\boldsymbol{x}_{i})\|^{2}$ is the regular data loss of PINN in matching the labled datasets; $\mathcal{L}_{gov}=\int_{\Omega}\|\mathcal{F}(u_{w};\theta)(\boldsymbol{x})\|^{2}\mathrm{d}\boldsymbol{x}$ is the residual loss that makes the network $u_{w}$ satisfy the governing equation constraints; $\mathcal{L}_{IC}=\int_{\Omega_{0}}\|\mathcal{I}(u_{w};\theta)(\boldsymbol{x})\|^{2}\mathrm{d}\boldsymbol{x}$ and $\mathcal{L}_{BC}=\int_{\partial\Omega}\|\mathcal{B}(u_{w};\theta)(\boldsymbol{x})\|^{2}\mathrm{d}\boldsymbol{x}$ are the corresponding losses that make PINN satisfy the initial and boundary conditions, respectively.
@@
In order to compute (4), several integral terms need to be evaluated, which involve the computation of higher-order derivatives of $u_{w}(\boldsymbol{x})$. PINN utilizes automatic differentiation of computational maps to calculate these derivative terms. The integrals are then approximated using a set of collocation points, which can be sampled using the Monte-Carlo method. We use $N_{d}$, $N_{g}$, $N_{i}$, $N_{b}$ to denote the number of corresponding data or collocation points. Then, the loss function can be approximated as [33]
$\begin{split}\mathcal{L}_{total}=\frac{\lambda_{d}}{N_{d}}\sum_{i=1}^{N}\|u_{w}(\boldsymbol{x}_{i})-u(\boldsymbol{x}_{i})\|^{2}+\frac{\lambda_{g}}{N_{g}}\sum_{i=1}^{N_{g}}\|\mathcal{F}(u_{w};\theta)(\boldsymbol{x}_{i})\|^{2}+\\ \frac{\lambda_{i}}{N_{i}}\sum_{i=1}^{N_{i}}\|\mathcal{I}(u_{w};\theta)(\boldsymbol{x}_{i})\|^{2}+\frac{\lambda_{b}}{N_{b}}\sum_{i=1}^{N_{b}}\|\mathcal{B}(u_{w};\theta)(\boldsymbol{x}_{i})\|^{2}+\cdots.\end{split}$ (5)
@@ 2.1.3 Implementation Based on DeepXDE
We implemented soft-constrained PINN with DeepXDE [25] and compared the results with a PyTorch-based implementation [26] (see  2.2.2 for details). DeepXDE is well-structured and highly configurable. Code written in DeepXDE is shorter, closer to mathematical formulas, and more computationally efficient. Solving differential equations in DeepXDE is akin to “building blocks” using built-in modules that specify the computational domain (geometric and temporal), differential equations, boundary/initial conditions, constraints, training datasets, neural network structure, optimization algorithms, and hyper-parameters, etc., all of which are loosely coupled.
@@ 2.2 Results of Soft-constrained PINN Solving Oscillator ODEs
For direct comparison, we choose to solve the same linear and nonlinear oscillator ODEs (i.e., Tutorial Example and Van der Pol osillators) as in [26].
@@ 2.2.1 PINN vs NN to Solving Linear Oscillator ODE
The only difference between solving with the soft-constrained PINN approach in this example and a conventional NN is the addition of $\mathcal{L}_{phys}$ to the total loss function. Since the initial condition in the conventional NN is already included in the training set (i.e., the first training point on the left), only $\mathcal{L}_{gov}$ satisfying Eq. (6) is added to total loss. This is accomplished by adding extra collocation points (see Eq. (5) ), in this example sampling at least 50 collocation points uniformly throughout the entire time domain. It could be understood that the training points are supervised and the collocation points are unsupervised. The exact number of collocation points required depends on the number of neural network parameters, i.e. the size of the problem being solved. The distribution of collocation points (uniform/random) also has an impact on the results [27]. The learning rate and loss weights also affect the convergence of the gradient descent algorithm [26]. Here the loss weights were taken as $\lambda_{d}=1.0,\lambda_{r}=6\times 10^{-2}$ through hyperparameter optimization. The loss weights act as a balance between the dual drivers of data and physical information.
@@
Figure 2: Result Comparison of PINN vs NN Solving Linear Oscillator–Case 1. $N_{data}$–No. of training data, $N_{c}$–No. of collocation points, all the same below. (Reproduced from [26] via DeepXDE for comparison.)
@@
Figure 3: Result Comparison of PINN vs NN Solving Linear Oscillator–Case 2. (Reproduced from [26] via DeepXDE.)
[…]
Minimalistic Training Data So, what is the minimum amount of training data required for the PINN method? For this example, if only one point, the initial value, is taken and the collocation points are kept to be 50, a good solution is obtained, as shown in Fig. 4(a), and the loss function descent process is shown in Fig. 4(b). [26] has already showed that when the nonlinearity of the problem is weak, a very small amount of labeled training data plus a few collocation points is sufficient to predict the solution. In the minimalist case, just as classical analytic or numerical integration methods require only one and two initial conditions for solving first- and second-order differential equations, respectively. PINN also require only one training point (initial value) and two training points containing the initial value point, respectively (see Fig. 7(b)).
@@
Figure 4: Minimalistic Training Data Example of First-order Linear ODE. (Reproduced from [26] via DeepXDE.)
[…]
Accelerating Training Time Compared to NN, PINN can significantly shorten the training time because it requires only a small amount of data (minimal labeled training data plus a small number of collocation points). On the other hand, as its loss function becomes more complex with the addition of $\mathcal{L}_{phys}$, it requires more computations and sometimes even more iteration epoches when training, which in turn drags down the training efficiency. We performed a validation test of Primer oscillator on DeepXDE 1.11.1 based on a CPU platform (Windows10, Intel Core i9-9900K @3.6GHz, 32GB Memory). The training time is the average of 5 training sessions (24,000 epoches per session, usually 24,000 epoches are not needed before convergence, and 24,000 epoches are trained for fair comparison), and the results are shown in Fig. 5. It can be seen that, for linear oscillator ODEs, the training is accelerated when the gains obtained along with the reduction in the amount of data outweigh the delay caused by the increase in the loss function terms. It should be emphasized that since Primer oscillator is a simple linear ODE, we used the same base network ($3\times 32$) for ease of comparison. PINN can accelerate the training even further by reducing the number of network parameters, because PINN need less network parameters than NN due to its small data characteristics. For example, for Fig. 5(b), if the base network is reduced to $3\times 16$, it also converges well while the training time is reduced to 15.30s.
@@
Figure 5: Training Time of PINN vs NN Solving Primer Oscillator via DeepXDE: (a) 16.56s by NN with Min. 50 training points; (b) 15.23s by PINN with Min. 1 initial value training point + 50 collocation points. (Reproduced from [26] via DeepXDE.)
@@ 2.2.2 PINN Solving Non-linear Oscillator ODE
The soft-constrained PINN approach also performs well in solving nonlinear oscillator ODEs [26]. We successfully solved the typical strongly nonlinear Van der Pol (VDP) oscillator based on DeepXDE, and further demonstrated the tolerance of PINN to noisy data as well as the light-code and training speed advantages of the DeepXDE implementation of PINN. The ODE of VDP is
@@
[…]
Noisy Data Tolerated The PINN method is used to solve the nonlinear VDP oscillator with small and noisy data. As with solving the Primer oscillator, the fully connected base network of $3\times 32$ was still used, which reflects the strong expressive power of neural networks inherited by PINN. The relevant parameter settings (i.e., $\epsilon$, number of training data points, number of collocation points, normal noise variance) are illustrated in Fig. 6, and other hyperparameter settings can be found in the code. Exact solution was generated by the 4th-order Runge-Kutta method. As can be seen from Fig. 6: 1) With the same amount of training data and collocation points, the results obtained from training with noisy data (right column) are basically identical to the noiseless results (left column), showing the good inclusiveness of the PINN method against noise. 2) As the degree of nonlinearity increases (by increasing $\epsilon$), more training and collocation points are needed to be arranged in regions with significant nonlinear features to obtain an exact solution. When $\epsilon=1,3,5$, the least number of training and collocation points are $N_{data}=28,32,38;N_{c}=15,25,40$, respectively. And 3) as the nonlinearity increases, the tolerance to noise decreases: when $\epsilon$ increases from 1 to 5, the $\sigma$ that can be tolerated decreases from 0.1 to 0.05; even increasing the number of data points has a limited effect on improving noise tolerance.
@@
[…]
PINN Implementation Code via DeepXDE vs PyTorch Comparison of Code  1 and Code  2 shows that the overall process (seen in Fig. 1) of realizing soft-constrained PINN based on DeepXDE or PyTorch is consistent. DeepXDE has a better customized encapsulation of the PINN method and thus requires less coding. With the same setup for solving VDP, the number of code lines (about 40 lines) via DeepXDE implementation is roughly one-third as the PyTorch implementation. In addition, the training efficiency of DeepXDE is significantly better than that of PyTorch, thanks to the optimization of DeepXDE for PINN [25]. For the example in Figure 5(b), PyTorch-based training took 43.10s with the same settings, which is about 1.8 times more than that of DeepXDE.
@@ 48 Loss_total = Loss_data + Loss_phys
[…]
PINN Computational Efficiency @CPU vs @GPU The PINN method does not depend on big data so it can be efficiently computed on platforms such as CPU or GPU. We compared the training time for solving Primer, VDP oscillators based on DeepXDE on CPU and GPU (CPU as before, GPU is Nvidia GTX1080Ti) respectively, taking the average of 5 training sessions (24000 epoches each session). Base networks use $2\times 32$ and $3\times 32$ fully connected hidden layers, respectively. As shown in Table  2, the training acceleration effect of GPU on PINN solving Primer and VDP is obvious: the training duration is reduced from 16.02s and 17.60s to 8.73s and 13.15s, respectively. When the base network is reduced to $3\times 16$, the training duration results show a consistent trend.
Table 2: PINN Training Time (Second) @CPU vs @GPU.
PINN to ODEs via DeepXDE @CPU @GPU